A manometer has a wall cross-sectional area of 4 cm3 and a tube cross-sectional area of 2 cm3 the tube is inclined at an angle 30 degree celsius to the well. If the well is exposed to a source of pressure, which result in the drop of the liquid of 10cm3. Calculate;1. The height of liquid dropped in well. 2. The height of liquid raised in the tube.
1
Expert's answer
2021-07-16T08:43:13-0400

Given,

Cross-sectional area of the wall of the manometer "(A_1)=4cm^2"

Cross-sectional area of the wall of the tube "(A_1)=2cm^2"

Angle of inclination of the tube "(\theta)=30^\circ"

Force applied by the source in the side of well = Force applied by the water due to raised in the height of liquid

"\Rightarrow" "P\times \Delta A\sin(30^\circ)=\rho \times V\times g"


"\Rightarrow P = \frac{\rho\times V\times g}{\Delta A \sin(30^\circ)}"


Now, substituting the values,


"\Rightarrow P= \frac{1000\times 10 \times 10^{-6}\times 10}{(4-2)\times 10^{-4}}N/m^2"


"\Rightarrow P = \frac{10^{3}}{2}N/m^2"

"\Rightarrow P =500Pa"